### v1

@@ Abstract
Results. The modest target densities for hadronic processes and the overall gamma-ray orbital variability favor inverse Compton scattering of ultraviolet photons from the massive companion star by highly relativistic electrons. The acceleration of the highest energy electrons implies a constraint on synchrotron cooling in the acceleration region, which can set an upper limit on its magnetic field. Moreover, the detected variability requires very strong particle confinement in both the acceleration and emission regions, which sets a lower limit on their magnetic fields that is barely consistent with the synchrotron cooling constraint from acceleration. Synchrotron losses may be higher in the emitting region if it is separated from the accelerator, but this requires a very hard particle injection spectrum.
@@ 1 Introduction
Twenty years ago, the High Energy Stereoscopic System (H.E.S.S.) detected orbitally modulated, very-high-energy (VHE; $100\,{\rm GeV}\lesssim\epsilon\lesssim 100$ TeV) emission and photons with energies reaching $\epsilon\sim 30$ TeV from LS 5039 (Aharonian et al., 2006a), a high-mass binary considered to be a likely GeV-emitting microquasar at the time (Paredes et al., 2000). Some years later, Fermi confirmed the source to be a GeV emitter (Abdo et al., 2009; Hadasch et al., 2012), and a re-analysis of MeV data strengthened previous evidence of detection in the MeV range (Collmar & Zhang, 2014). This information was complemented by soft and hard X-ray observations, which found an orbital light curve similar to that in TeV (e.g., Bosch-Ramon et al., 2005; Hoffmann et al., 2009; Takahashi et al., 2009), and to that hinted by the MeV data re-analysis. The High-Altitude Water Cherenkov Observatory (HAWC) has recently detected the source up to $\epsilon\approx 200$ TeV, reaching the photon ultra-high-energy (UHE; $\epsilon>100$ TeV) range, with a spectral energy distribution (SED) $\propto\epsilon^{-0.8}$ that shows no evidence of a high-energy cut-off (Alfaro et al., 2025). HAWC has also found $2.7\sigma$ evidence of orbital variability in the $40-118$ TeV range, and $2.2\sigma$ evidence above 100 TeV. Furthermore, the emission detected around the inferior conjunction of the compact object (INFC; orbital phase range $0.45-0.9$) reaches 208 TeV, whereas around the superior conjunction (SUPC) the emission reaches 118 TeV, which also suggests variability above 100 TeV.
@@
LS 5039 was initially considered to be a microquasar due to the presence of X-ray accretion features and radio jets (Ribó et al., 1999; Paredes et al., 2000). However, several authors proposed that this system may host a young pulsar instead (Martocchia et al., 2005; Dubus, 2006b), as in the case of PSR B1259-63/LS2883, which is also a TeV-emitting binary (Johnston et al., 1992; Aharonian et al., 2005). In addition, it was found that the previously found X-ray accretion features were likely to be background contamination due to the poor angular resolution of RXTE (Ribó et al., 1999; Bosch-Ramon et al., 2005). However, while a pulsar-star wind collision scenario is generally accepted to be behind the nonthermal (NT) emission in PSR B1259-63/LS2883 (e.g., Tavani & Arons 1997; Khangulyan et al. 2007; see however Yi & Cheng 2017), in LS 5039 a microquasar scenario cannot yet be fully discarded. The compact object mass has not been ascertained (Casares et al., 2005), the presence of pulsations is controversial (Yoneda et al., 2020; Kargaltsev et al., 2023; Makishima et al., 2023), and some accretion modes may not yield prominent features such as thermal emission and spectroscopic lines (Okazaki et al., 2008; Barkov & Khangulyan, 2012). More exotic scenarios have been discussed as well, such as a young magnetar, prompted by the hint of pulsations with a period of $\approx 9$ s found in X-rays (Yoneda et al., 2020). Such an object, however, seems to be at odds with the absence of a conspicuous supernova remnant surrounding the source (Ribó et al., 2002; Moldón et al., 2012), and assessing its feasibility needs further theoretical development (Bosch-Ramon & Barkov, in prep.). Therefore, the debate about the processes yielding the NT emission in LS 5039 remains far from being solved, a debate that is tightly connected to other similar systems. In particular, in LS I+61 303, which presents evidence for 269 ms radio pulsations (Weng et al., 2022) and thus disfavors the microquasar scenario (see however Jaron et al., 2024), the question of the origin of the NT emission is still open (e.g., a colliding-wind model vs. a transitional – magnetar-like – pulsar scenario; e.g., Dubus, 2006b; Papitto et al., 2012).
@@
In this work, we aimed to improve our knowledge of the NT accelerator and emitter in LS 5039 by exploring the consequences of HAWC observational results using analytical tools. This analysis, approximate but suitable for this source due to the large uncertainties involved, is based on the spectral and variability properties of the UHE emission and is conducted within the framework of IC scattering off stellar photons. We took $\epsilon=100$ TeV as the reference photon energy to study the orbital variability, a compromise between assuming that the modulation is only real up to $\epsilon\sim 40$ TeV, or that it actually goes beyond $\sim 100$ TeV as INFC photons reach $\approx 208$ TeV while SUPC photons reach just $\approx 118$ TeV, and Alfaro et al. 2025 found $2.2\sigma$ evidence of orbital variability above 100 TeV. This work strengthens the conclusions of Khangulyan et al. (2008), gets deeper into the global implications for the emitter and the accelerator, and proposes an acceleration mechanism based on Derishev et al. (2003) and Derishev & Aharonian (2012). The paper is organized as follows. The emitting and the acceleration regions are discussed in Sects. 2 and 3, respectively, and the suggested model is proposed in Sect. 4. We conclude with Sect. 5.
@@ 2 The UHE photon emitter
Anisotropic IC scattering and $\gamma\gamma$ absorption represent two geometrical factors that may influence the orbital phase dependence of the gamma-ray emission, but their influence is strongly suppressed in the UHE regime. Thus, the orbital UHE variability is caused by one or several of these other factors changing along the orbit: hydrodynamics (i.e., escape, adiabatic losses, or heating); intrinsic emission beaming (i.e., Doppler boosting or an anisotropic electron angular distribution); and IC target density. All these effects (including the minor geometric ones) can only be present if the emitter size ($R_{\rm em}$) is not significantly larger than $a$, as otherwise orbital variability would be washed out. Moreover, the mean free path, $\lambda,$ of the highest energy particles in the emitter, at least equal to their gyroradius $r_{\rm g}=E/q{\cal B}_{\rm em}$, where $q$ is the absolute value of the particle charge and ${\cal B}_{\mathrm{em}}$ is the magnetic field strength in the emitter, should fulfill
$\quantity(\frac{\lambda}{r_{\rm g}})r_{\rm g}\lesssim R_{\rm em}\lesssim a\,.$ (1)
Thus, electron confinement requires an emitter magnetic field
@@
Further constraints can be derived from the spectral slope in the UHE regime. Since for $120$ TeV electrons the KN parameter reaches $E\epsilon_{\star}/m^{2}_{e}c^{4}\sim 5\times 10^{3}$, one can adopt the asymptotic relation between the photon and the electron spectra. In particular, the gamma-ray differential distribution $dN_{\epsilon}/d\epsilon\propto\epsilon^{-2.83\pm 0.09^{+0.02}_{-0.05}}$ around INFC (Alfaro et al., 2025) requires an emitting electron spectrum similar to $dN/dE\propto E^{-2}$ within uncertainties. Synchrotron losses modify the electron injection spectrum ($dQ/dE=dN_{\rm inj}/dEdt$) as follows:
@@
where $v_{\mathrm{em}}$ is the emitting flow velocity and a spherical geometry is assumed. The equation was derived taking into account that the Poynting flux cannot overcome the total power of the source ($L_{\rm w}$), and that the latter minimum value is that required to feed the NT emission in LS 5039 (i.e., $L_{\mathrm{NT}}\gtrsim 10^{36}$ erg s-1 from the overall MeV-peaked SED; Collmar & Zhang 2014). If electrons escaped to the stellar wind region and radiated there, they would find a magnetic field:
@@ 4 Acceleration model
Synchrotron cooling and confinement decouple when particles that are roughly isotropic in the laboratory frame (LF) are immersed in a magnetized relativistic outflow. Under these conditions, the flow frame (FF) magnetic field is able to deflect the particles enough for them to acquire a large amount of energy in the LF (as demonstrated in Sects. 4.1 and 4.2), while the rate of synchrotron cooling in the FF is significantly reduced (as shown in Sect. 4.3). The presence of a bright stellar companion provides a way to bring relativistic particles into the relativistic outflow: LS 5039 is optically thick for gamma rays with energies between $\sim 10$ GeV and 10 TeV interacting with stellar photons on binary scales, so secondary $e^{\pm}$ pairs (secondaries hereafter) are expected to be injected within the binary (e.g., Dubus, 2006a; Aharonian et al., 2006b), and thus within the outflow produced by the compact object (Derishev & Aharonian, 2012). These secondaries, which constitute the seed particles for the acceleration process, are mostly created with energies in the range $\sim 10-100$ GeV, depending on their injection location in the system with respect to the gamma-ray emitter (Bosch-Ramon et al., 2008a). Secondaries could also be created in the local NT emitter photon fields [^2] Properly accounting for these local photon fields requires a detailed model of the overall NT emitter, which is beyond the scope of this work., although we focus on those from $\gamma\gamma$ absorption in the stellar photon field because they reach the highest energies once accelerated.
@@
where both $E^{\prime}$ and $r^{\prime}_{\rm g}$ are constant in the FF. Most of the secondaries injected with energy $E_{0}$, if moving isotropically in the LF, have similar though slightly lower $E^{\prime}$ values. For simplicity, we assumed no electric field in the wind FF, that is, the plasma has infinite conductivity. A sketch of the process from the point of view of both the FF and the LF is shown in Fig. 1.
Equation (16) provides an illustrative way to find an upper limit for the maximum change in the particle energy in the LF. If a particle gyrates for exactly half an orbit starting from an initial orientation against the wind, then its maximum vertical displacement is $2r^{\prime}_{\mathrm{g}}$ in the FF; in the LF, the maximum particle vertical displacement is $2r^{\prime}_{\mathrm{g}}$ as well. Since in the LF the electric field, ${\cal E}_{\mathrm{acc}}$ , is also perpendicular to the wind bulk speed, the maximum possible energy gain is thus
@@ 4.1 Maximum energy
$x=\Gamma\quantity(x^{\prime}(t^{\prime})+\beta ct^{\prime})\,.$ (21)
@@
For relatively small $E_{0}$ values (e.g., $\lesssim 10$ MeV for $\Gamma\sim 10^{3}$), secondaries gyrate up to $\theta^{\prime}\gtrsim\pi/2$, and Eq. (18) yields $E\sim\Gamma^{2}E_{0}$ (see Eq. 17 above). For the higher $E_{0}$ values given above though ($\sim 10-100$ GeV), $\theta^{\prime}$ can be $\ll 1$. In that case, we adopted a third order Taylor expansion:
@@
which in fact yields results correct within $\sim 10$% even for $\theta^{\prime}\sim 1$.
From Eqs. (21) and (23), the first order Taylor expansion of $\beta_{\mathrm{w}}\approx 1-1/2\Gamma^{2}$, the $t^{\prime}$-$\theta^{\prime}$ relation, and the fact that in the present case $1/(1-(v^{\prime}/c)^{2})^{1/2}\gg\Gamma$, we derived for $x(\theta^{\prime})$
@@
similar to those required to explain HAWC observations. Higher energies could be achieved for larger $E_{0}$ values, but the number of secondaries strongly diminishes well above $100$ GeV. We note that, for a final $\theta^{\prime}<\sqrt{3}/\Gamma$, $x$ in Eq. (24) is negative and $E$ barely grows above $E_{0}$. This applies in the unlikely event that $E_{0}$ approaches $E_{\rm H}$, so particles escape almost undeflected in the FF.
@@ 4.4 Electron spectrum and emitting region
i) The accelerator features a hard particle spectrum up to $E_{\rm max}$: While still in the wind zone, radiative losses remain negligible, but once particles are injected downstream of the wind termination shock, synchrotron losses render $dN/dE\propto E^{-2}$ for a $dQ/dE\propto E^{-1}$ or harder. The IC emission from these particles is consistent with the HAWC SED, although their corresponding synchrotron emission would not be consistent with the spectrum at lower energies requiring another emitting particle component (or region), as already discussed at the end of Sect. 2. This would be compatible with the fact that a significant part of the X-rays and VHE gamma rays seem to come from a region with a size $\gtrsim a$ (e.g., Bosch-Ramon et al., 2007; Khangulyan et al., 2008; Bosch-Ramon et al., 2008b; Szostek & Dubus, 2011).
@@ 5 Conclusions
Recent results from HAWC show that LS 5039 features a relatively hard SED and orbital variability up to UHE. Most of this emission, like the VHE emission, is produced in the INFC. This, together with the system properties, strongly suggests the same radiation mechanism in both energy bands, most likely IC scattering off stellar photons. The exceedingly high energies of the electrons behind the UHE emission, and the emission orbital modulation, imply very extreme conditions of the accelerator and the emitter. Even under the most conservative assumption that variability only affects photon energies up to $\approx 40$ TeV, synchrotron losses must be very small in the accelerator, but the corresponding magnetic field can hardly confine the particles in the region. A potential solution to this problem is a version of the converter mechanism proposed by Derishev & Aharonian (2012) in the context of a high-mass gamma-ray binary hosting an ultrarelativistic outflow. To obtain the maximum electron energies, we proposed a simplified model based on that mechanism that significantly alleviates the tensions, as synchrotron cooling is reduced without having to decrease the magnetic field. Detailed, thorough studies are nevertheless needed to properly understand the process quantitatively; this would better characterize the properties of the injected secondaries, the accelerated particle spectrum, the connection between the UHE and the broadband NT emitter, and the specific UHE variability cause(s).

### v2

@@ Abstract
Results. The modest target densities for hadronic processes and the overall gamma-ray orbital variability favor inverse Compton scattering of ultraviolet photons from the massive companion star by highly relativistic electrons. The acceleration of the highest energy electrons implies a constraint on synchrotron cooling in the acceleration region, which can set an upper limit on its magnetic field. Moreover, the detected variability requires very strong particle confinement in both the acceleration and emission regions, which sets a lower limit on their magnetic fields that is barely consistent with the synchrotron cooling constraint from acceleration. Synchrotron losses may be higher in the emitting region if it is separated from the accelerator, but this requires a very hard particle injection spectrum. An accelerator based on an ultrarelativistic magnetized outflow can alleviate these requirements.
@@ 1 Introduction
Twenty years ago, the High Energy Stereoscopic System (H.E.S.S.) detected orbitally modulated, very-high-energy (VHE; $100\,{\rm GeV}\lesssim\epsilon\lesssim 100$ TeV) emission and photons with energies reaching $\epsilon\sim 30$ TeV from LS 5039 (Aharonian et al., 2006a), a high-mass binary considered to be a likely GeV-emitting microquasar at the time (Paredes et al., 2000). Some years later, Fermi confirmed the source to be a GeV emitter (Abdo et al., 2009; Hadasch et al., 2012), and a re-analysis of MeV data strengthened previous evidence of detection in the MeV range (Collmar & Zhang, 2014). This information was complemented by soft and hard X-ray observations, which found an orbital light curve similar to that in TeV (e.g., Bosch-Ramon et al., 2005; Hoffmann et al., 2009; Takahashi et al., 2009), and to that hinted by the MeV data re-analysis. The High-Altitude Water Cherenkov Observatory (HAWC) has recently detected the source up to $\epsilon\approx 200$ TeV, reaching the photon ultra-high-energy (UHE; $\epsilon>100$ TeV) range, with a spectral energy distribution (SED) $\propto\epsilon^{-0.8}$ that shows no evidence of a high-energy cut-off (Alfaro et al., 2025). HAWC has also found $2.7\sigma$ evidence of orbital variability in the 4011840-11840 - 118 TeV range, and $2.2\sigma$ evidence above 100 TeV. Furthermore, the emission detected around the inferior conjunction of the compact object (INFC; orbital phase range 0.450.90.45-0.90.45 - 0.9) reaches 208 TeV, whereas around the superior conjunction (SUPC) the emission reaches 118 TeV, which also suggests variability above 100 TeV.
@@
LS 5039 was initially considered to be a microquasar due to the presence of X-ray accretion features and radio jets (Ribó et al., 1999; Paredes et al., 2000). However, several authors proposed that this system may host a young pulsar instead (Martocchia et al., 2005; Dubus, 2006b), as in the case of PSR B1259-63/LS2883, which is also a TeV-emitting binary (Johnston et al., 1992; Aharonian et al., 2005). In addition, it was found that the previously found X-ray accretion features were likely to be background contamination due to the poor angular resolution of RXTE (Ribó et al., 1999; Bosch-Ramon et al., 2005). However, while a pulsar-star wind collision scenario is generally accepted to be behind the nonthermal (NT) emission in PSR B1259-63/LS2883 (e.g., Tavani & Arons 1997; Khangulyan et al. 2007; see however Yi & Cheng 2017), in LS 5039 a microquasar scenario cannot yet be fully discarded. The compact object mass has not been ascertained (Casares et al., 2005), the presence of pulsations is controversial (Yoneda et al., 2020; Kargaltsev et al., 2023; Makishima et al., 2023), and some accretion modes may not yield prominent features such as thermal emission and spectroscopic lines (Okazaki et al., 2008; Barkov & Khangulyan, 2012). More exotic scenarios have been discussed as well, such as a young magnetar, prompted by the hint of pulsations with a period of 9$\approx$ 9≈ 9 s found in X-rays (Yoneda et al., 2020). Such an object, however, seems to be at odds with the absence of a conspicuous supernova remnant surrounding the source (Ribó et al., 2002; Moldón et al., 2012), and assessing its feasibility needs further theoretical development (Bosch-Ramon & Barkov, in prep.). Therefore, the debate about the processes yielding the NT emission in LS 5039 remains far from being solved, a debate that is tightly connected to other similar systems. In particular, in LS I +61 303, which presents evidence for 269 ms radio pulsations (Weng et al., 2022) and thus disfavors the microquasar scenario (see however Jaron et al., 2024), the question of the origin of the NT emission is still open (e.g., a colliding-wind model vs. a transitional – magnetar-like – pulsar scenario; e.g., Dubus, 2006b; Papitto et al., 2012).
@@
In this work, we aimed to improve our knowledge of the NT accelerator and emitter in LS 5039 by exploring the consequences of HAWC observational results using analytical tools. This analysis, approximate but suitable for this source due to the large uncertainties involved, is based on the spectral and variability properties of the UHE emission and is conducted within the framework of IC scattering off stellar photons. We took $\epsilon=100$ TeV as the reference photon energy to study the orbital variability, a compromise between assuming that the modulation is only real up to $\epsilon\sim 40$ TeV, or that it actually goes beyond 100$\sim$ 100∼ 100 TeV as INFC photons reach 208$\approx$ 208≈ 208 TeV while SUPC photons reach just 118$\approx$ 118≈ 118 TeV, and Alfaro et al. 2025 found $2.2\sigma$ evidence of orbital variability above 100 TeV. The present work strengthens the conclusions of Khangulyan et al. (2008), gets deeper into the global implications for the emitter and the accelerator, and proposes an acceleration mechanism based on Derishev et al. (2003) and Derishev & Aharonian (2012). The paper is organized as follows. The emitting and the acceleration regions are discussed in Sects. 2 and 3, respectively, and the suggested model is proposed in Sect. 4. We conclude with Sect. 5.
@@ 2 The UHE photon emitter
Anisotropic IC scattering and $\gamma\gamma$ absorption represent two geometrical factors that may influence the orbital phase dependence of the gamma-ray emission, but their influence is strongly suppressed in the UHE regime. Thus, the orbital UHE variability is caused by one or several of these other factors changing along the orbit: hydrodynamics (i.e., escape, adiabatic losses, or heating); intrinsic emission beaming (i.e., Doppler boosting or an anisotropic electron angular distribution); and IC target density. All these effects (including the minor geometric ones) can only be present if the emitter size ($R_{\rm em}$) is not significantly larger than $a$, as otherwise orbital variability would be washed out. Moreover, the mean free path, $\lambda,$ of the highest energy particles in the emitter (at least equal to their gyroradius $r_{\rm g}=E/q{\cal B}_{\rm em}$, where $q$ is the absolute value of the particle charge and ${\cal B}_{\mathrm{em}}$ is the magnetic field strength in the emitter) should fulfill $(\lambda/r_{\rm g})r_{\rm g}\lesssim R_{\rm em}\lesssim a$. Thus, electron confinement requires an emitter magnetic field
@@
Further constraints can be derived from the spectral slope in the UHE regime. Since for 120120120 TeV electrons the KN parameter reaches $E\epsilon_{\star}/m^{2}_{e}c^{4}\sim 5\times 10^{3}$, one can adopt the asymptotic relation between the photon and the electron spectra. In particular, the gamma-ray differential distribution $dN_{\epsilon}/d\epsilon\propto\epsilon^{-2.83\pm 0.09^{+0.02}_{-0.05}}$ around INFC (Alfaro et al., 2025) requires an emitting electron spectrum similar to $dN/dE\propto E^{-2}$ within uncertainties. Synchrotron losses modify the electron injection spectrum ($dQ/dE=dN_{\rm inj}/dEdt$) as follows:
@@
where $v_{\mathrm{em}}$ is the emitting flow velocity and a spherical geometry is assumed. The equation was derived taking into account that the Poynting flux cannot overcome the total power of the source ($L_{\rm w}$), and that this power minimum value is that required to feed the NT emission in LS 5039 (i.e., $L_{\mathrm{NT}}\gtrsim 10^{36}$ erg s-1 from the overall MeV-peaked SED; Collmar & Zhang 2014). If electrons escaped to the stellar wind region and radiated there, they would find a magnetic field:
@@ 4 Acceleration model
Synchrotron cooling and confinement decouple when particles that are roughly isotropic in the laboratory frame (LF) are immersed in a magnetized relativistic outflow. Under these conditions, the flow frame (FF) magnetic field is able to deflect the particles enough for them to acquire a large amount of energy in the LF (as demonstrated in Sects. 4.1 and 4.2), while the rate of synchrotron cooling in the FF is significantly reduced (as shown in Sect. 4.3). The presence of a bright stellar companion provides a way to bring relativistic particles into the relativistic outflow: LS 5039 is optically thick for gamma rays with energies between 10$\sim$ 10∼ 10 GeV and 10 TeV interacting with stellar photons on binary scales, so secondary $e^{\pm}$ pairs (secondaries hereafter) are expected to be injected within the binary (e.g., Dubus, 2006a; Aharonian et al., 2006b), and thus within the outflow produced by the compact object (Derishev & Aharonian, 2012). These secondaries, which constitute the seed particles for the acceleration process, are mostly created with energies in the range 10100$\sim$ 10-100∼ 10 - 100 GeV, depending on their injection location in the system with respect to the gamma-ray emitter (Bosch-Ramon et al., 2008a). Secondaries could also be created in the local NT emitter photon fields [^2] Properly accounting for these local photon fields requires a detailed model of the overall NT emitter, which is beyond the scope of this work., although we focus on those from $\gamma\gamma$ absorption in the stellar photon field because they reach the highest energies once accelerated.
@@
where both $E^{\prime}$ and $r^{\prime}_{\rm g}$ are constant in the FF in a planar wind. Most of the secondaries injected with energy $E_{0}$, if moving isotropically in the LF, have similar though slightly lower $E^{\prime}$ values. For simplicity, we assumed no electric field in the wind FF, that is, the plasma has infinite conductivity. A sketch of the process from the point of view of both the FF and the LF is shown in Fig. 1.
[…]
@@ 4.1 Maximum energy
$x=\Gamma\quantity(x^{\prime}(t^{\prime})+\beta_{\rm w}ct^{\prime})\,.$ (20)
@@
For relatively small $E_{0}$ values (e.g., 10$\lesssim$ 10≲ 10 MeV for $\Gamma\sim 10^{3}$), secondaries gyrate up to $\theta^{\prime}\gtrsim\pi/2$, and Eq. (17) yields $E\sim\Gamma^{2}E_{0}$ (see Eq. 16 above). For the higher $E_{0}$ values given above though (10100$\sim$ 10-100∼ 10 - 100 GeV), $\theta^{\prime}$ can be 1$\ll$ 1≪ 1. In that case, we adopted a third order Taylor expansion:
@@
which in fact yields results correct within 10$\sim$ 10∼ 10% even for $\theta^{\prime}\sim 1$.
From Eqs. (20), (21) and (22), the first order Taylor expansion of $\beta_{\mathrm{w}}\approx 1-1/2\Gamma^{2}$, the $t^{\prime}$-$\theta^{\prime}$ relation, and the fact that in the present case $1/(1-(v^{\prime}/c)^{2})^{1/2}\gg\Gamma$, we derived for $x(\theta^{\prime})$
@@
similar to those required to explain HAWC observations. Higher energies could be achieved for larger $E_{0}$ values, but the number of secondaries strongly diminishes well above 100100100 GeV. We note that, for a final $\theta^{\prime}<\sqrt{3}/\Gamma$, $x$ in Eq. (23) is negative and $E$ barely grows above $E_{0}$. This applies in the unlikely event that $E_{0}$ approaches $E_{\rm H}$, so particles escape almost undeflected in the FF.
@@ 4.4 Electron spectrum and emitting region
i) The accelerator features a hard particle spectrum up to $E_{\rm max}$: While still in the wind zone, radiative losses remain negligible, but once particles are injected downstream of the wind termination shock, synchrotron losses render $dN/dE\propto E^{-2}$ for a $dQ/dE\propto E^{-1}$ or harder. The IC emission from these particles is consistent with the HAWC SED, although their corresponding synchrotron emission would not be consistent with the spectrum at lower energies requiring another emitting particle component (or region), as already discussed at the end of Sect. 2. This would be compatible with the fact that a significant part of the X-rays and VHE gamma rays seem to come from a region with a size close to but $\gtrsim a$ (e.g., Bosch-Ramon et al., 2007; Khangulyan et al., 2008; Bosch-Ramon et al., 2008b; Szostek & Dubus, 2011).
@@ 5 Conclusions
Recent results from HAWC show that LS 5039 features a relatively hard SED and orbital variability up to UHE. Most of this emission, like the VHE emission, is produced in the INFC. This, together with the system properties, strongly suggests the same radiation mechanism in both energy bands, most likely IC scattering off stellar photons. The exceedingly high energies of the electrons behind the UHE emission, and the emission orbital modulation, imply very extreme conditions of the accelerator and the emitter. Even under the most conservative assumption that variability only affects photon energies up to 40$\approx$ 40≈ 40 TeV, synchrotron losses must be very small in the accelerator, but the corresponding magnetic field can hardly confine the particles in the region. A potential solution to this problem is a version of the converter mechanism proposed by Derishev & Aharonian (2012) in the context of a high-mass gamma-ray binary hosting an ultrarelativistic outflow. To obtain the maximum electron energies, we proposed a simplified model based on that mechanism that significantly alleviates the tensions, as synchrotron cooling is reduced without having to decrease the magnetic field. Detailed, thorough studies are nevertheless needed to properly understand the process quantitatively; this would better characterize the properties of the injected secondaries, the accelerated particle spectrum, the connection between the UHE and the broadband NT emitter, and the specific UHE variability cause(s).